OpenStudy (anonymous):

Find the exact area of the region bounded by two concentric circles with radii 10 inches and 6 inches. i think its 64 pi correct me if I'm wrong?

OpenStudy (amistre64):

might need some context ....

OpenStudy (amistre64):

concentric ... common centers, thats what i was missing

OpenStudy (amistre64):

area big - area little

OpenStudy (amistre64):

pi R^2 - pir^2 = pi(R^2-r^2)

OpenStudy (anonymous):

so.. is my answer wrong or right? and that the question i got.. im confuzed by it too

OpenStudy (amistre64):

100 - 36 = 64 yes so 64pi

OpenStudy (amistre64):

64pi "in^2"

OpenStudy (anonymous):

yay i got one right for once :D thank you!

OpenStudy (amistre64):

youre welcome; i had to remember what concentric circles were .. but after that it was simple enough